OpenStudy (anonymous):

Prove that \(\lim_{N\rightarrow \infty}N\sum_{n=N}^{\infty}\frac{1}{n^2}\) exists.

OpenStudy (xapproachesinfinity):

why prove question draw attention lol

OpenStudy (xapproachesinfinity):

questions*

geerky42 (geerky42):

Because it is interesting question

OpenStudy (anonymous):

my first guess, which may be wrong, is to prove that this is cauchy sequence

OpenStudy (anonymous):

Since \(N\rightarrow\infty\), N>1. Moreover, \(\frac{1}{n^2}\ge0 \) for all \(n\in N\). Hence, we have \(0\le\sum_{n=N}^{\infty}\frac{1}{n^2}\le\sum_{n=1}^{\infty}\frac{1}{n^2}\). Thus, \(0\le N\sum_{n=N}^{\infty}\frac{1}{n^2}\le N\sum_{n=1}^{\infty}\frac{1}{n^2}\). But then, I am not sure how I can proceed from here, or if I am on the right track...

OpenStudy (xapproachesinfinity):

@geerky42, I agree lol

OpenStudy (anonymous):

If it is a Cauchy sequence, then the series converges. But what's next?

OpenStudy (anonymous):

i have no idea how to show it is cauchy, that was just my guess as a strategy

OpenStudy (anonymous):

Does "it" refer to \(N\sum_{n=N}^{\infty}\frac{1}{n^2}\) or \(\frac{1}{n^2}\) ?

OpenStudy (anonymous):

if you are asking me i would say "it" refers to \(N\sum_{n=N}^{\infty}\frac{1}{n^2}\)

OpenStudy (zarkon):

you are working too hard

OpenStudy (anonymous):

oh damn

OpenStudy (zarkon):

try an integral test

OpenStudy (anonymous):

don't tell me it is l'hopital or something that simple

OpenStudy (anonymous):

We haven't learnt it in our class yet D:

OpenStudy (zarkon):

you haven't done the integral test?

OpenStudy (anonymous):

Not yet. Only Cauchy, Comparison, Ratio, Root, and Alternating Series Test.

OpenStudy (zarkon):

well an integral test is derived from the comparison test

OpenStudy (zarkon):

\[\sum_{n=N}^\infty\frac{1}{n^2}\le \int\limits_{N}^{\infty}\frac{1}{x^2}dx =\frac{1}{N}\]

OpenStudy (anonymous):

yeah, it is one damn

OpenStudy (zarkon):

\[N\sum_{n=N}^\infty\frac{1}{n^2}\le N\int\limits_{N}^{\infty}\frac{1}{x^2}dx =N\frac{1}{N}=1\]

OpenStudy (zarkon):

not real happy though...still missing something

OpenStudy (zarkon):

need to show that \(\displaystyle N\sum_{n=N}^\infty\frac{1}{n^2}\) as a sequence in N is monotone

OpenStudy (zarkon):

it is monotone. I can show the difference \(a_{N+1}-a_N\) is negative

OpenStudy (zarkon):

bounded monotone sequences converge

OpenStudy (zarkon):

i should have \[N\sum_{n=N}^\infty\frac{1}{n^2}\le N\int\limits_{N-1}^{\infty}\frac{1}{x^2}dx =N\frac{1}{N-1}=\frac{1}{1-\frac{1}{N}}\]

OpenStudy (anonymous):

Not happy at all... \[a_{N+1}-a_{N} = (N+1)\sum_{n=N+1}^{\infty}\frac{1}{n^2}-N\sum_{n=N}^{\infty}\frac{1}{n^2} = N(\sum_{n=N+1}^{\infty}\frac{1}{n^2} - \sum_{n=N}^{\infty}\frac{1}{n^2}) + \sum_{n=N+1}^{\infty}\frac{1}{n^2}\]\[=-\frac{1}{N}+\sum_{n=N+1}^{\infty}\frac{1}{n^2}\] Well...

OpenStudy (zarkon):

\[1= N\int\limits_{N}^{\infty}\frac{1}{x^2}dx \] \[\le N\sum_{n=N}^\infty\frac{1}{n^2}\le N\int\limits_{N-1}^{\infty}\frac{1}{x^2}dx =N\frac{1}{N-1}=\frac{1}{1-\frac{1}{N}}\] by the squeze theorem \[1\le \lim_{N\to\infty}N\sum_{n=N}^\infty\frac{1}{n^2}\le\lim_{N\to\infty}\frac{1}{1-\frac{1}{N}}=1\] the limit is 1

OpenStudy (zarkon):

using what I have above you don't need to show the sequence is monotone.

OpenStudy (zarkon):

so not only does it converge we know the limit now

OpenStudy (anonymous):

We are not expected to use integral test though

OpenStudy (zarkon):

I didn't technically. I used the comparison test

OpenStudy (zarkon):

the integral test just tells you if a series converse or not. It doesn't give you bounds